OpenStudy (anonymous):

Write balanced chemical equations for the following: a. A solution of sulfuric acid (battery acid) is neutralized with a solution of sodium hydroxide. b. The combustion of iron metal produces iron(III) oxide (rust). c. When sodium metal is dropped into water, it fizzes. d. The science fair “volcano reaction” is produced by mixing sodium bicarbonate (baking soda) with vinegar (a solution of acetic acid HC2H3O2).

OpenStudy (anonymous):

I believe this is the answer to A. H_2^+ 〖(SO〗_4^(2-) )_2+2〖Na〗^+ O^(2-) H^+→〖2Na〗^+ 〖SO〗_4^(2-)+H_4^+ O_2^(2-)

OpenStudy (jebonna):

A. 2 NaOH + H2SO4 = Na2SO4 + 2 H2O B. 4 Fe + 3 O2 = 2 Fe2O3 C. 2Na + 2H20 = 2NaOH + H2 D. NaHCO3 + HC2H3O2 = CO2 + H2O + NaC2H3O2